# CALRec: Contrastive Alignment of Generative LLMs For Sequential Recommendation

### v1

@@ 3.5. Quasi-Round-Robin BM25 Retrieval
At inference, we prompt the model with the text description of user’s history of interactions. To obtain more than one candidate next item prediction, we conduct temperature sampling (Hinton et al., 2015) to generate $N_{\text{gen}}$ pieces of text as $N_{\text{gen}}$ predictions for the next item. Each of the $N_{\text{gen}}$ predictions is associated with its sequence score $s$ (i.e., log probability, which is a negative real number typically in the range of $(-10,0)$, as observed on our data).
@@
Input : $samples$: Output texts sampled from the LLM (size = $N_{\text{gen}}$ )
@@
Input : $logprobs$: LLM prediction scores for $samples$ (size = $N_{\text{gen}}$)
@@
Input : $C$: Item corpus (size = $N_{c}$)
Input : $N_{\text{preds}}$: a hyperparam controlling the number of generated text samples to use for BM25 matching.
Input : $\epsilon$: a hyperparam used to scale the LLM log probs score before combining with the BM25 score.
Output : The sorted list of item predictions.
4
@@
6
7sorted__\__samples, sorted__\__logprobs $\leftarrow$ RemoveDuplicates((sorted__\__samples, sorted__\__logprobs))
@@
       // Find BM25 match scores of entire corpus for each samples
@@
       // Modulate by sequence prediction score.
@@
We remove duplicate generations, sort them by sequence scores in descend order, and retain at most top $N_{\text{preds}}$ unique predictions. Then, in order to match these generated texts with items from our corpus, we rely on BM25 retrieval (Robertson et al., 2009) to calculate the matching scores between each text prediction and the whole item corpus (with size $N_{c}$): this will result in BM25 matching scores organized into a matrix of shape ($N_{c}$, $N_{\text{preds}}$). In order to rank the candidate items, we devise the following heuristics which aims to select the closest matches according to both the LLM scores and BM25 scores in a quasi round robin fashion along the different $N_{\text{preds}}$ text generations. First, we linearly scale each column of the matrix respectively to the range $[0,1]$ to make BM25 scores derived from different LLM predictions comparable. We then penalize each prediction’s BM25 scores (each column) by multiplying the BM25 scores by a factor of $e^{\epsilon s}$, where $s$ is the sequence score associated with the text prediction, and $\epsilon$ is a small positive constant so that $e^{\epsilon s}$ is smaller than but very close to $1.0$. This is to further modulate the BM25 text-based retrieval scores by the LLM generative scores. After that, we max-pool along the $N_{\text{preds}}$ axis and finally derive modulated BM25 matching scores in a single vector of size $N_{c}$, with information from all $N_{\text{preds}}$ recommendations fused in it (we can then derive the ranking of the ground-truth target item in the item corpus for calculating evaluation metric scores, or return top-K recommendations for users), as presented in Algorithm 1222In our preliminary investigation, we found that 1) max-pooling significantly outperforms min-pooling and mean-pooling and 2) there is obvious performance drop if we remove modulation since only max-pooling combined with modulation can make sure that top-1 LLM outputs’ top-1 recommendation is still the top-1 recommendation in the merged list. 3) The overall results are not sensitive to the value of $\epsilon$ as long as $\epsilon$ is small.. Here, we provide another view into the round-robin nature of our approach: given that the scaled BM25 score of each best-match item from the non-duplicate $N_{\text{preds}}$ texts generated is $1$, when they are further modulated (scaled down) using $e^{\epsilon s}$, the final retrieval algorithm roughly follows a round-robin fashion, where the top-1 BM25 retrieval result of each of the generated text are in turn grouped as the final top $N_{\text{preds}}$ retrieval results, when $\epsilon$ is close to $0.0$. Since there are ties, we name our method ’quasi-’ rather than ’strict-’ round-robin.
@@ 4. Experimental Setup
Table 1. Statistics of Amazon Reviews Dataset.
LLM Selection. Our work leverages PaLM-2 (Anil et al., 2023; Tay et al., 2023), a state-of-the-art LLM pretrained with a mixture of different objectives and demonstrates superb generalization capabilities on a variety of language tasks. Among PaLM-2 variations, we choose PaLM-2 XXS as the backbone for our tasks for trade-off in model capability and inference latency. Larger PaLM-2 size-S does not show a significant advantage against XXS without fine-tuning (see §5). Besides, the scale of our dataset (significantly smaller than text corpora for pretraining LLMs) may lead to overfitting when finetuning larger models (Kirkpatrick et al., 2016; Mitchell et al., 2022).
@@
User Sequence Deduplication. We have found that the Amazon Review datasets include a non-negligible percentage of consecutive duplicate events in user sequences, where the duplication is identified as two consecutive activities by the same user that have the same item ID (ASIN number), review text, rating, and the same review timestamp. As indicated in Table 1, around $4\%$ to $9.4\%$ of purchased items are exact duplicates of the previous ones in their user sequences. We conduct deduplication (dedup) on all user sequences following our strict definition of duplication above, while we still keep two consecutive purchases of the same item, if they occurred at different time (i.e., they come with different timestamps) or come with different ratings/review text since they are likely the users’ real actions. We note that some previous work did not conduct user sequence deduplication (Li et al., 2023b), which may lead a recommender system to learn a trivial strategy of always recommending the last item in the input sequence.
@@
Implementation Details. We train our CALRec based on XXS-sized PaLM-2 on a cluster of TPUv4 chips (Jouppi et al., 2023); our code is implemented with JAX (Bradbury et al., 2018) [^4] https://github.com/google/jaxand PAX [^5] https://github.com/google/paxml. We adopt a training batch size of $512$ users, a learning rate of $1$e44-4- 4, a maximum input length of $1,024$ tokens, and a maximum output decoding length of $80$ tokens. As a data augmentation trick, during training, we randomly truncate the last $k$ items ($k\in\{0,1,2,3,4\}$) from the full user record (with $n$ items) and always use the last item after truncation (i.e., the $(n-k)$-th item in the original user record) as the target item. We train CALRec with $\alpha=0.125,\beta=-0.025$ for up to $150,000$ steps in Stage I and adopt $\tau_{c}=0.5$ (see Eqs. 23) as the config combination yields best validation scores. We adopt the checkpoint with $135,000$ training steps in Stage I selected on the validation data category, and then further fine-tune the model for Stage II. There are two exceptions: Scientific [^6] We pick the Stage I model at $45,000$ training steps since ‘Scientific’ is significantly smaller than other categories and its own validation set indicates that overfitting occurs at $135,000$ steps and then continue with Stage II training., and Games [^7] For Games category, we adopt $\alpha=0.125,\beta=0$ for both stages and its Stage I checkpoint at $105,000$ steps is selected using the Stage I validation set. It is because we found on the ‘Games’ own validation set, the new setup outperforms the previous setup . All our model selection is based on the NDCG@10 metric following previous work (specifically we adopt its pessimistic estimation as defined in the last paragraph of this section). For each case, the same training objective is used in Stages I and II. At inference, the temperature for decoding is $0.5$, $N_{gen}=32$, $N_{pred}=10$ and $\epsilon=1/5000$. There are two hyper-parameters inherent to BM25: we adopt the default values from the original BM25 API, $k_{1}=1.5$ and $b=0.75$, throughout all our experiments and did not further tune them since we found that the model performance is not sensitive to them.
@@
Baselines. The following are all the baseline approaches we considered. SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019) are pure ID-Based approaches, Recformer (Li et al., 2023b) and the ‘Last Item Repeater’ (LIR) are purely text-based, and FDSA (Zhang et al., 2019) and S3-Rec (Zhou et al., 2020) leverage both ID and text. UniSRec (Hou et al., 2022) has two variants and we report its performance in both text-only mode and ID+text mode. More details about each baseline are provided in what follows.  [^8] In our experiments, LIR is implemented on our own, Recformer follows its official implementation. The rest baselines are all provided by the official code repo of UniSRec which is based on RecBole (Zhao et al., 2021): the attribute embeddings of FDSA and S3-Rec are text features from our strongest baseline UniSRec.
@@
  • FDSA (Zhang et al., 2019) applies two self-attentional blocks to address ID sequence and attribute feature sequence, respectively, concatenates their final representations and fuses them for next item prediction.
@@
  • UniSRec (Hou et al., 2022) separately embeds each item’s text via a pretrained and fixed BERT cascaded by a trainable adapter. Given a sequence of item embeddings, a sequence representation is then modeled with another small transformer network. We adopt the officially released UniSRec model pretrained on multiple domains and fine-tune it on specific Amazon categories, respectively. We report two fine-tuning setups of UniSRec: (1) text-only setup, (2) ID+text setup where an item embedding is the sum of its text and ID representations.
@@
  • Recformer (Li et al., 2023b) relies on the encoder-only Longformer architecture (Beltagy et al., 2020), token type and item position embeddings are proposed to facilitate learning item and user representations. It also features a pretraining stage on multiple categories with a combination of standard Masked Language Modeling (MLM) loss and a contrastive objective aligning user history and target item representations. We adopt Recformer’s officially released pretrained checkpoint and fine-tune it on specific categories, respectively.
@@
  • Last Item Repeater, or LIR hereafter, denotes a simple yet effective baseline strategy which always recommends the last item in the input sequence as our training-free, text-statistic baseline. Based on the text of the last item, we rely on BM25 to rank the items in the full item collection.
@@
Evaluation Metrics. Following standard practices from prior work, we report NDCG@10, Recall@K (K$\in\{1,10\}$) and MRR@$\infty$ scores. Different from previous work, we report both ‘optimistic’ (opt) and ‘pessimistic’ (pes) calculations of the metrics: when there is a tie such that the ground-truth target item and other candidate items have the same matching score, the ground-truth item ranks first in (opt) and ranks last in (pes) calculation, respectively. It is because 1) there are distinct items (with different item IDs) with the same text description and 2) in rare cases, we found that BM25 may yield identical scores for very similar text pieces. We still treat items in case 1) as distinct items because in real production environments vendors/content creators may intentionally or unintentionally duplicate items. Of course, this will not cause any issue for ID-based approaches due to distinct item IDs, where (opt) and (pes) scores are the same. In the official implementations of our text-based baselines, Li et al. (2023b) only calculates (opt) metric scores which may result in unfair comparisons with ID-based baselines.  Hou et al. (2022) applies “torch.topk()” to derive exactly top $K$ item ids for calculating Recall/NDCG@K: if the $K$-th and the $K+1$-th items are having the same score, only one of them will be included in the output list. So its metric score is a number between our(opt) and (pes) estimates. [^9] To provide more insights on our (opt)/(pes) metrics, readers could consider an extreme scenario where the model output is empty and thus when using BM25 retrieval the matching score is 00 for all candidate items. In this case, the optimistic Recall@K scores will always be $1.0$, while the pessimistic scores will always be $0.0$. To provide a fair assessment of the model’s ability to uniquely identify candidate items, we have thus decided to report both (opt) and (pes) scores. All the evaluations are performed on user-sequence-deduplicated datasets.
@@ 5. Experimental Results
Table 2. Evaluation results of CALRec in comparison with different ID/Text-based sequential recommendation baselines (cf. §4). For text-only methods, we report both optimisitc (opt) and pessimistic (pes) as “(opt)/(pes)”. For ID-based methods, (opt) and (pes) scores are the same. Bold: the highest (opt) scores. Underline: the strongest (pes) scores. We also report the relative improvement (opt) comparing CALRec against (opt) baseline SotA marked with Blue Background per data category per metric.
@@
- Multi-Category Joint Fine-Tuning (A) 0.0727 / 0.0675 0.04408 / 0.0314 0.1070 / 0.1061 0.0670 / 0.0602 0.0870 / 0.0862 0.0673 / 0.0658 0.1119 / 0.1115 0.0822 / 0.0812 0.0583 / 0.0570 0.0286 / 0.0265 0.0966 / 0.0959 0.0501 / 0.0486
- Category-Specific Fine-Tuning (B) 0.0717 / 0.0670 0.0424 / 0.0300 0.1077 / 0.1073 0.0653 / 0.0588 0.0867 / 0.0862 0.0715 / 0.0702 0.1048 / 0.1048 0.0840 / 0.0834 0.0539 / 0.0530 0.0257 / 0.0238 0.0904 / 0.0902 0.0463 / 0.0451
- Contrastive Alignment 0.0771 / 0.0724 0.0495 / 0.0376 0.1109 / 0.1105 0.0712 / 0.0649 0.0905 / 0.0899 0.0711 / 0.0696 0.1154 / 0.1153 0.0858 / 0.0850 0.0593 / 0.0584 0.0289 / 0.0272 0.0987 / 0.0985 0.0507 / 0.0495
@@
Table 3. Ablation result showing the benefits of multi-stage finetuning and constrastive alignment to improve the performance of CALRec. Bold: the highest optimistic (opt) scores. Underline: the strongest pessimistic (pes) scores.
Metric W/ Contr. W/O Contr. Improv. p-value
@@
Pretrained PaLM-2 W/O Fine-Tuning
@@
Table 5. Performance results of pretrained PaLM-2 without any fine-tuning for Sequential Recommendation. Bold: the highest optimistic (opt) scores. Underline: the strongest pessimistic (pes) scores.
@@
Main Results. Table 2 demonstrates that the sequential recommendation capabilities of the proposed CALRec framework are stronger than our baselines. On the $6$ Amazon Review test categories, our model outperforms all $8$ baselines significantly in terms of NDCG@10, Recall@1 and MRR. In terms of Recall@10, our approach outperforms all Text-Only and ID-Only models across the board by a large margin, but slightly underperforms our strongest baseline the ID+Text variant of UniSRec. For our method and $3$ Text-Only baselines, the results in Table 2 verified that there is a gap between optimistic and pessimistic metric scores, demonstrating the necessity of reporting both. Meanwhile, it is worth noticing that our pessimistic scores clearly surpass the optimistic scores of the $3$ text-only models on all $6$ data categories. In addition to the absolute scores, we also calculate the percentage gains comparing CALRec’s optimistic scores against the baseline SotA per each data category and per evaluation metric (different method for different data category and metric). The results are presented in the last column of Table 2. We found that, CALRec is especially strong in Recall@1, outperforming baseline SotA by $37.3\%$ on average; CALRec outperforms baseline SotA by $23.8\%$ and $22.3\%$ on average in terms of NDCG@10 and MRR, respectively.
@@
Ablation Study. Finetuning Stages: In Table 3, we demonstrate the effectiveness of CALRec’s key components via extensive ablation study on $3$ data categories (the results on the other $3$ categories show the same trends so we skip them due to space constraints). First, it is obvious that our Stage I multi-category joint fine-tuning and Stage II category-specific fine-tuning are both effective, without which there will be an obvious drop in performance. Furthermore, it is worth mentioning that our model derived only with Stage I is able to recommend items in all $6$ categories: this is especially valuable and easy to deploy in real-world applications, due to its multi-tenancy nature. Although the multi-category trained model (variant B in Table 3) lags behind the full CALRec derived with both stages, it is already on par with variant A models (cf. Table 3)). We also compare CALRec and pretrained ‘off-the-shelf’ PaLM-2 (XXS) in Table 3; the results show that without any fine-tuning, PaLM-2 demonstrates poor performance on our sequential recommendation benchmarks.
@@
Contrastive Objectives: In Table 4 we calculate the average scores on all $6$ categories for the setups with and without the auxiliary contrastive losses training objective. Basically, the auxiliary contrastive objective can result in circa $0.8\%-1.7\%$ gains. We also conduct dependent paired sample t-test (Ross and Willson, 2017) and report the p-values in the last column of Table 4: the gains are all statistically significant at $0.05$ level.
@@
Model Size: We are also interested in how pretrained PaLM-2 without any fine-tuning on our sequential recommendation data performs in our tasks and we present results with three PaLM-2 models of three different model sizes (i.e., XXXS, XXS, and S) on ‘Scientific’ and ‘Instruments’ categories in Table 5. The input/output format is still the same: input data is from a single user. The results show that XXS and S models actually achieve similar performance and they both outperform the XXXS model.

### v2

@@ 3.5. Quasi-Round-Robin BM25 Retrieval
At inference, we prompt the model with the text description of user’s history of interactions (see also §3.2). To obtain more than one candidate next-item prediction, we conduct temperature sampling (Fan et al., 2018) to generate $N_{\text{gen}}$ pieces of text as $N_{\text{gen}}$ predictions for the next item. Each of the $N_{\text{gen}}$ predictions is associated with its sequence score $s$ (i.e., log probability, which is a negative real number typically in the range of $(-10,0)$, as observed on our data).
@@
Input :  $samples$, output texts sampled from the LLM (size = $N_{\text{gen}}$);
$logprobs$, LLM sequence scores of $samples$ (size = $N_{\text{gen}}$);
$C$, corpus of candidate items (size = $N_{c}$);
$N_{\text{preds}}$, a hyperparameter that determines the number of generated
text samples used for BM25 matching;
$\epsilon$, a hyperparameter that controls the modulation of BM25 scores
with LLM sequence scores.
@@
Output : The sorted list of item predictions.
@@
5
6sorted__\__samples, sorted__\__logprobs $\leftarrow$ RemoveDuplicates((sorted__\__samples, sorted__\__logprobs))
7
// Keep the top $N_{\text{preds}}$ samples
@@
       // Find BM25 scores between entire corpus and each sample
@@
       // Modulate by sequence score
@@
We remove duplicate generations, sort them by sequence scores in descend order, and retain at most top $N_{\text{preds}}$ unique predictions. Then, in order to match these generated texts with items from our corpus, we rely on BM25 retrieval (Robertson et al., 2009) to calculate the matching scores between each text prediction and the whole item corpus (with size $N_{c}$): this will result in BM25 matching scores organized into a matrix of shape ($N_{c}$, $N_{\text{preds}}$). In order to rank the candidate items, we devise the following heuristics which aims to select the closest matches according to both the LLM sequence scores and BM25 scores in a quasi-round-robin fashion across the different $N_{\text{preds}}$ text generations. First, we linearly scale each column of the matrix respectively to the range $[0,1]$ to make BM25 scores derived from different LLM predictions comparable. We then penalize each prediction’s BM25 scores (each column) by multiplying the BM25 scores by a factor of $e^{\epsilon s}$, where $s$ is the sequence score associated with the text prediction, and $\epsilon$ is a small positive constant so that $e^{\epsilon s}$ is smaller than but very close to $1.0$. This is to further modulate the BM25 text-based retrieval scores by the LLM generative scores. After that, we max-pool along the $N_{\text{preds}}$ axis and finally derive modulated BM25 matching scores in a single vector of size $N_{c}$, with information from all $N_{\text{preds}}$ recommendations fused in it (we can then derive the ranking of the ground-truth target item in the item corpus for calculating evaluation metric scores, or return top-K recommendations for users), as presented in Algorithm 1333In our preliminary investigation, we found that 1) max-pooling significantly outperforms min-pooling and mean-pooling and 2) there is obvious performance drop if we remove modulation since only max-pooling combined with modulation can make sure that top-1 LLM outputs’ top-1 recommendation is still the top-1 recommendation in the merged list. 3) The overall results are not sensitive to the value of $\epsilon$ as long as $\epsilon$ is small.. Here, we provide another view into the round-robin nature of our approach: given that the scaled BM25 score of each best-match item from the non-duplicate $N_{\text{preds}}$ texts generated is $1$, when they are further modulated (scaled down) using $e^{\epsilon s}$, the final retrieval algorithm roughly follows a round-robin fashion, where the top-1 BM25 retrieval result of each of the generated text are in turn grouped as the final top $N_{\text{preds}}$ retrieval results, when $\epsilon$ is close to $0.0$. Since there are ties, we name our method ‘quasi-’ rather than ‘strict-’ round-robin.
@@ 4. Experimental Setup
LLM Selection. Our work leverages PaLM-2 (Anil et al., 2023; Tay et al., 2023), a state-of-the-art LLM that is pretrained with a mixture of different objectives and exhibits exceptional generalization capabilities on a variety of language tasks. Among PaLM-2 model variants, we choose PaLM-2 XXS as the LLM backbone for our tasks, balancing model capability with inference latency. [^4] Here, XXS is a ‘T-shirt size’ notation of the model size. S-, M- and L-sized PaLM-2 models reported in Anil et al. (2023) are all larger than our XXS-sized model. Larger PaLM-2 S does not show a significant advantage against XXS model when both are without fine-tuning (see §5). Besides, the scale of our dataset (significantly smaller than text corpora for pretraining LLMs) may lead to overfitting when fine-tuning larger models (Kirkpatrick et al., 2017; Mitchell et al., 2022).
@@
Table 1. Statistics of Amazon Review dataset.
@@
User Sequence Deduplication. We have found that the Amazon Review data include a non-negligible percentage of consecutive duplicate events in user sequences, where the duplication is identified as two consecutive activities by the same user that have the same item ID (ASIN number), review text, rating, and the same review timestamp. As indicated in Table 1, around $4\%$ to $9.4\%$ of purchases are exact duplicates of the previous ones in their user sequences. We conduct deduplication (dedup) on all user sequences following our strict definition of duplication above, while we still keep two consecutive purchases of the same item, if they occurred at different times (i.e., they come with different timestamps) or come with different ratings/review text since they are likely the users’ real actions. We note that some previous work, e.g.  (Li et al., 2023c), did not conduct user sequence deduplication, which may lead a recommender system to learn a trivial strategy of always recommending the last item in the input sequence.
@@
Implementation Details. We train our CALRec based on XXS-sized PaLM-2 on a cluster of TPUv4 chips (Jouppi et al., 2023); our code is implemented with JAX (Bradbury et al., 2018) [^7] https://github.com/google/jaxand PAX [^8] https://github.com/google/paxml. We adopt a training batch size of $512$ users, a learning rate of $1$e44-4- 4, a maximum input length of $1,024$ tokens, and a maximum output decoding length of $80$ tokens. As a data augmentation trick, during training, we randomly truncate the last $k$ items ($k\in\{0,1,2,3,4\}$) from the full user record (with $n$ items) and always use the last item after truncation (i.e., the $(n-k)$-th item in the original user record) as the target item. We train CALRec with $\alpha=0.125,\beta=-0.025$ for up to $150,000$ steps in Stage I and adopt $\tau_{c}=0.5$ (see Eqs. 23) as the configuration yields best validation scores. We adopt the checkpoint with $135,000$ training steps in Stage I selected on the validation data category, and then further fine-tune the model for Stage II. There are two exceptions: Scientific [^9] We pick the Stage I model at $45,000$ training steps since ‘Scientific’ is significantly smaller than other categories and its own validation set indicates that overfitting occurs at $135,000$ steps and then continue with Stage II training., and Games [^10] For Games category, we adopt $\alpha=0.125,\beta=0$ for both stages and its Stage I checkpoint at $105,000$ steps is selected using the Stage I validation set. It is because we found on the ‘Games’ own validation set, the new setup outperforms the previous (default) setup.. All our model selection is based on the NDCG@10 metric following previous work (specifically we adopt its pessimistic estimation which will be defined later in this section). For each case, the same training objective is used in Stages I and II. At inference, the temperature for decoding is $0.5$, $N_{gen}=32$, $N_{pred}=10$ and $\epsilon=1/5000$. There are two hyper-parameters inherent to BM25: we adopt the default values from the original BM25 API, $k_{1}=1.5$ and $b=0.75$, throughout all our experiments and did not further tune them since we found that the model performance is not sensitive to them.
@@
Baselines. In this work, we consider a comprehensive set of baseline approaches. SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019) are pure ID-Based approaches, Recformer (Li et al., 2023c) and the ‘Last Item Repeater’ (LIR) are purely text-based, and FDSA (Zhang et al., 2019) and S3-Rec (Zhou et al., 2020) leverage both ID and text. UniSRec (Hou et al., 2022) has two variants and we report its performance in both text-only mode and ID+text mode. More details about each baseline are provided in what follows. [^11] In our experiments, LIR is implemented on our own, Recformer follows its official implementation. The rest baselines are all provided by the official code repo of UniSRec which is based on RecBole (Zhao et al., 2021): the attribute embeddings of FDSA and S3-Rec are text features from our strongest baseline UniSRec.
@@
  • FDSA (Zhang et al., 2019) applies two self-attentional blocks to address ID sequence and attribute feature sequence, respectively, and fuses their final representations for next-item prediction.
@@
  • UniSRec (Hou et al., 2022) separately encodes each item’s text via a pretrained and fixed BERT cascaded by a trainable adapter. Given a sequence of item embeddings, a sequence representation is then modeled with another small transformer network. We adopt the officially released UniSRec model pretrained on multiple domains and fine-tune it on specific Amazon categories, respectively. We report two fine-tuning setups of UniSRec: (1) text-only setup, (2) ID+text setup where an item embedding is the sum of its ID and text representations.
@@
  • Recformer (Li et al., 2023c) relies on the encoder-only Longformer architecture (Beltagy et al., 2020) where token type and item position embeddings are proposed to facilitate learning item and user representations. It also features a pretraining stage on multiple categories with a combination of standard Masked Language Modeling (MLM) loss and a contrastive objective aligning user history and target item representations. We adopt Recformer’s officially released pretrained checkpoint and fine-tune it on specific categories, respectively.
@@
  • Last Item Repeater, or LIR hereafter, denotes a simple yet effective baseline strategy which always recommends the last item in the input sequence as our training-free, text-statistic baseline. Based on the text of the last item, we rely on BM25 to rank the items in the full item corpus.
@@
Evaluation Metrics. Following standard practices from prior work, we report NDCG@10, Recall@K (K$\in\{1,10\}$) and MRR@$\infty$ scores. Different from previous work, we report both ‘optimistic’ (opt) and ‘pessimistic’ (pes) calculations of the metrics: when there is a tie such that the ground-truth target item and other candidate items have the same matching score, the ground-truth item ranks first in (opt) and ranks last in (pes) calculation, respectively. It is because 1) there are distinct items (with different item IDs) with the same text description and 2) in rare cases, we found that BM25 may yield identical scores for very similar text pieces. We still treat items in case 1) as distinct items because in real production environments vendors/content creators may intentionally or unintentionally duplicate items. Of course, this will not cause any issue for ID-based approaches due to distinct item IDs, where (opt) and (pes) scores are the same. In the official implementations of our text-based baselines, Li et al. (2023c) only calculates (opt) metric scores which may result in unfair comparisons with ID-based baselines.  Hou et al. (2022) applies “torch.topk()” to derive exactly top $K$ item ids for calculating Recall/NDCG@K: if the $K$-th and the $K+1$-th items are having the same score, only one of them will be included in the output list. So its metric score is a number between our (opt) and (pes) estimates. [^12] To provide more insights on our (opt)/(pes) metrics, readers could consider an extreme scenario where the model output is empty and thus when using BM25 retrieval the matching score is 00 for all candidate items. In this case, the optimistic Recall@K scores will always be $1.0$, while the pessimistic scores will always be $0.0$. To provide a fair assessment of the model’s ability to uniquely identify candidate items, we have thus decided to report both (opt) and (pes) scores. All our evaluations are performed on user-sequence-deduplicated data.
@@ 5. Experimental Results
Table 2. Evaluation results of CALRec in comparison with different ID/Text-based sequential recommendation baselines (cf. §4). For text-only methods, we report both optimisitc (opt) and pessimistic (pes) as ‘(opt)/(pes)’. For ID-based methods, (opt) and (pes) scores are the same. Bold: the highest (opt) scores. Underline: the strongest (pes) scores. We also report the relative improvement (opt) comparing CALRec against (opt) baseline SotA marked with Blue Background per data category per metric. Evaluation metrics are computed using the Amazon Review dataset after user sequence deduplication (see also §4).
@@
w/o Multi-Category Joint Fine-Tuning (A) 0.0727 / 0.0675 0.04408 / 0.0314 0.1070 / 0.1061 0.0670 / 0.0602 0.0870 / 0.0862 0.0673 / 0.0658 0.1119 / 0.1115 0.0822 / 0.0812 0.0583 / 0.0570 0.0286 / 0.0265 0.0966 / 0.0959 0.0501 / 0.0486
w/o Category-Specific Fine-Tuning (B) 0.0717 / 0.0670 0.0424 / 0.0300 0.1077 / 0.1073 0.0653 / 0.0588 0.0867 / 0.0862 0.0715 / 0.0702 0.1048 / 0.1048 0.0840 / 0.0834 0.0539 / 0.0530 0.0257 / 0.0238 0.0904 / 0.0902 0.0463 / 0.0451
w/o Contrastive Alignment 0.0771 / 0.0724 0.0495 / 0.0376 0.1109 / 0.1105 0.0712 / 0.0649 0.0905 / 0.0899 0.0711 / 0.0696 0.1154 / 0.1153 0.0858 / 0.0850 0.0593 / 0.0584 0.0289 / 0.0272 0.0987 / 0.0985 0.0507 / 0.0495
@@
Table 3. Ablation result showing the benefits of multi-stage fine-tuning and constrastive alignment to improve the performance of CALRec. Bold: the highest optimistic (opt) scores. Underline: the strongest pessimistic (pes) scores.
Metric: (opt)/(pes) w/ Contr. w/o Contr. Improv. $p$-value
@@
Model Size
@@
Table 5. Performance results of pretrained PaLM-2 models of three distinct sizes without any fine-tuning for Sequential Recommendation. Bold: the highest optimistic (opt) scores. Underline: the strongest pessimistic (pes) scores.
Dataset Metric (opt)/(pes)
Scientific NDCG@10 0.0265/0.0232
Recall@1 0.0051/0.0030
Recall@10 0.0582/0.0537
MRR 0.0217/0.0189
Instruments NDCG@10 0.0240/0.0237
Recall@1 0.0028/0.0025
Recall@10 0.0584/ 0.0583
MRR 0.0179/0.0175
Table 6. Performance results of GPT4Rec’s template using pretrained PaLM-2 XXS without any fine-tuning for Sequential Recommendation.
@@
Main Results. Table 2 demonstrates that the sequential recommendation capabilities of the proposed CALRec framework are stronger than our baselines. On the $6$ Amazon Review test categories, our model outperforms all $8$ baselines significantly in terms of NDCG@10, Recall@1 and MRR. In terms of Recall@10, our approach outperforms all Text-Only and ID-Only models across the board by a large margin (including the Text-Only variant of UniSRec), but slightly underperforms our strongest baseline, the ID+Text variant of UniSRec. For our method and $3$ Text-Only baselines, the results in Table 2 verified that there is a gap between optimistic and pessimistic metric scores, demonstrating the necessity of reporting both. Meanwhile, it is worth noticing that even our pessimistic scores clearly surpass the optimistic scores of the $3$ Text-Only baseline models on all the $6$ data categories. In addition to the absolute scores, we also calculate the percentage gains comparing CALRec’s optimistic scores against the baseline SotA for each data category and each evaluation metric (different method for different data category and metric). The results are presented in the last column of Table 2. We found that, CALRec is especially strong in Recall@1, outperforming baseline SotA by $37.3\%$ on average; CALRec outperforms baseline SotA by $23.8\%$ and $22.3\%$ on average in terms of NDCG@10 and MRR, respectively.
Ablation Study. In Table 3, we demonstrate the effectiveness of CALRec’s key components via extensive ablation study on $3$ data categories (the results on the other $3$ categories show the same trends so we skip them due to space constraints).
@@
Two-Stage Fine-Tuning: First, it is obvious that our Stage I multi-category joint fine-tuning and Stage II category-specific fine-tuning are both effective, without which there will be an obvious drop in performance. Furthermore, it is worth mentioning that our model derived only with Stage I is able to recommend items in all $6$ categories: this is especially valuable and easy to deploy in real-world applications, due to its multi-tenancy nature. Although the multi-category trained model (variant B in Table 3) lags behind the full CALRec derived with both stages, it is already on par with variant A models (cf. Table 3). We also compare CALRec and pretrained ‘off-the-shelf’ PaLM-2 XXS in Table 3; the results show that without any fine-tuning, PaLM-2 demonstrates poor performance on our sequential recommendation benchmarks.
@@
Contrastive Objectives: In Table 4 we calculate the average scores on all $6$ categories for the setups with and without the auxiliary contrastive losses. Basically, the auxiliary contrastive objective can result in circa $0.8\%-1.7\%$ gains. We also conduct dependent paired sample t-test (Ross and Willson, 2017) and report the $p$-values in the last column of Table 4: the gains are all statistically significant at $0.05$ level.
@@
Model Size: We are also interested in how pretrained PaLM-2 without any fine-tuning on our sequential recommendation data performs in our tasks and we present results with three PaLM-2 models of different model sizes (i.e., XXXS, XXS, and S) on ‘Scientific’ and ‘Instruments’ categories in Table 5. The input/output format is still the same: input data is from a single user. The results show that XXS and S models actually achieve similar performance and they both outperform the XXXS model.
Template Design: We conducted additional experiments on prompting the raw PaLM2 XXS model for recommendation adopting another prompt from GPT4Rec (Li et al., 2023d), with results reported in Table 6. Unlike our prompt template, their template doesn’t include a shared per-item prefix and user suffix but leverages a one-off user suffix. Compared with our results reported in Table 5 with the same raw PaLM-2 XXS, the new results are universally lower than ours (especially recall@1, relatively $30\%\sim 45\%$ drop in scores), showing the effectiveness of our template design.